### v1

# Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA) for Automatic Alignment in Large Language Models

Abstract
Recent advancements in Large Language Model (LLM) alignment have sought to mitigate the cost of human annotations by leveraging pretrained models to generate preference data. However, existing methods often compare responses from models with substantially different capabilities, yielding superficial distinctions that fail to provide meaningful guidance on what constitutes a superior response. To address this limitation, we propose Kinship-Aware pReference MApping (KARMA), a novel framework that systematically pairs responses from models with comparable competencies. By constraining preference comparisons to outputs of similar complexity and quality, KARMA enhances the informativeness of preference data and improves the granularity of alignment signals. Empirical evaluations demonstrate that our kinship-aware approach leads to more consistent and interpretable alignment outcomes, ultimately facilitating a more principled and reliable pathway for aligning LLM behavior with human preferences.
Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA)
for Automatic Alignment in Large Language Models
Jeesu Jung1, Chanjun Park2, Sangkeun Jung1††thanks: Corresponding author 1Chungnam National University, 2Korea University jisu.jung5@gmail.com, bcj1210@korea.ac.kr, hugmanskj@gmail.com
1 Introduction
Aligning Large Language Models (LLMs) with human preferences is a fundamental challenge in artificial intelligence, where the effectiveness of alignment hinges on the quality and specificity of the preference data used to guide model outputs Shen et al. . Early alignment methodologies primarily relied on human-annotated datasets with binary preference labels, wherein explicit human judgments directed models toward generating more desirable responses Ouyang et al. (2022a). While these methods have played a critical role in improving LLM behavior, their dependence on extensive human labor imposes significant scalability constraints.
To address this limitation, recent research has explored automated preference data generation techniques, such as Reinforcement Learning from AI Feedback (RLAIF) Lee et al. . These methods leverage pretrained models to automatically generate preference pairs, significantly reducing the need for human annotation. However, the quality of the resulting preference data remains a major concern. If responses within a preference pair exhibit substantial disparities in quality, complexity, or style, the preference signal may become trivial, capturing only superficial contrasts rather than informative distinctions that refine model behavior.
We define this issue as the Response Quality Gap, which refers to the disparity in coherence, complexity, and reasoning quality between candidate responses. A large Response Quality Gap weakens preference signals, leading to misaligned or less informative training objectives. This problem is particularly pronounced in large-scale, automatically generated datasets, where manually curating response pairs to minimize the Response Quality Gap is infeasible due to the massive volume of data.
A promising solution lies in considering the origin of the responses being compared—specifically, the LLM that generated them. Since a model’s reasoning ability, commonsense knowledge, and instruction-following capabilities are inherently shaped by its training data and architecture, the quality of its responses is largely determined by these characteristics. Consequently, comparing responses from models with vastly different capabilities introduces inconsistencies in preference data, making alignment objectives less reliable.
To formalize this idea, we introduce the notion of Kinship among LLMs, which quantifies the similarity in reasoning ability, commonsense understanding, and instruction adherence among models. Measured using standard benchmarks, kinship provides a principled way to assess whether two models generate responses of comparable complexity and quality. Building on this insight, we propose the Kinship-Aware pReference MApping (KARMA) framework, which systematically pairs responses based on model kinship. Unlike conventional approaches that indiscriminately compare outputs from models of varying capabilities, KARMA ensures that preference pairs are drawn from models with similar competencies, thereby narrowing the Response Quality Gap. This results in preference data that captures more subtle and meaningful distinctions, leading to higher-quality alignment signals.
The effectiveness of KARMA is illustrated in Figure 1, which provides an overview of the framework, and Figure 2, which details its response evaluation process. By fostering more structured and granular comparisons, KARMA improves the coherence and interpretability of preference data, shifting alignment objectives from filtering out trivial differences to refining preference signals through comparability and consistency. Our empirical results demonstrate that this kinship-aware approach significantly outperforms prior methods such as RLAIF, offering a more robust and scalable solution for aligning LLM behavior with human preferences.
2 Kinship-Aware pReference MApping (KARMA) Framework
Recent approaches to LLM alignment have reduced the reliance on human annotation through automated preference data generation. However, these methods often neglect the foundational influence of the Origin Model. For example, a model trained using Llama’s methodology is going to produce outputs that fall within the distribution of Llama’s training data. Just as a son inherits traits from his father, a model’s quality directly influences the quality of its outputs. when preference pairs are derived from models with markedly disparate capabilities, the resulting comparisons yield only trivial distinctions.
To address this, the proposed Kinship-Aware pReference MApping (KARMA) framework evaluates kinship based on the performance of origin models on selected benchmark metrics. With kinship as a guiding principle for structuring binarized data, the framework systematically generates preference data for alignment tuning.
2.1 Definition of Kinship
In the context of binarized alignment data mapping, the similarity between the responses serves as a fundamental determinant of mapping efficacy. When data pairs are closely aligned, they provide more nuanced and meaningful guidance signals for alignment tuning. Building upon this principle of response similarity, KARMA introduces the concept of kinship among models.
The kinship is measured by evaluating the capabilities of origin models. We measure the benchmarks as the origin model capability. We define a set of candidate origin models, , each evaluated according to multiple dimensions of competency. Let be a set of benchmark tasks. Each benchmark provides an evaluation metric that quantifies the performance of model . This evaluation may include measures of factual correctness, reasoning, and adherence to instructions.
For each pair of models , we define a kinship score , which reflects the similarity of their capabilities:
where is a similarity function. Higher values of indicate closer competency alignment between and . The exact functional form of can vary depending on the chosen benchmarks and may incorporate multiple factors (denoted collectively as , , and ) to minimize bias and ensure comprehensive evaluation.
Descriptive Multiple-Choice Instruction Following IF evalZhou et al. (2023) Knowledge MMLUHendrycks et al. (2020) MMLU-proWang et al. (2024) Reasoning ARC-EasyClark et al. (2018) ARC-ChallengeClark et al. (2018) HellaswagZellers et al. (2019)
2.2 Stage 1: Benchmark Evaluation for Model Competency
To identify which models are suitable candidates for kinship-based pairing, we employ a series of benchmarks that meet three criteria: (1) they are widely validated in prior research, (2) they cover both general knowledge and domain-specific tasks, and (3) they provide well-defined quantitative metrics. A summary of selected benchmarks is provided in Table 1.
Evaluating models on these benchmarks yields performance scores that serve as the basis for computing . The degree of kinship among models varies with the choice of benchmarks. To avoid relying solely on log-probabilities or overly simplistic metrics, we adopt prompt-based evaluations that require each model to produce explicit responses. Appendix C presents the benchmark performance of all models used for kinship calculation, along with example calculations.
The prompt formats are adapted from templates in MMLU-pro Wang et al. (2024) and FLAN Longpre et al. (2023), ensuring compatibility with each model’s input structure and recommended usage patterns. The prompt template can be found in Appendix B.
2.3 Stage 2: Kinship-Based Response Mapping
Once we identify a subset of models whose pairwise kinship scores exceed a predetermined threshold (e.g., ), we set at 0.1, as cosine similarity values below this threshold indicate a lack of meaningful similarity between models. we focus on collecting responses from these closely matched models. Kinship serves as an indicator of model’s ability and similarity. While various criteria can be applied to measure this, we utilize the models’ capabilities and performance similarities that can be estimated through benchmarks. We have defined kinship across three distinct cases:
: Mapping the two origin models with the highest average benchmark scores.
: Mapping the two origin models with the highest cosine similarity between their benchmark scores.
: Mapping the model with the highest benchmark score to the most similar model based on cosine similarity.
Figure 3 shows the each kinship definition and calculation process.
To generate preference data from these kinship relationships, we convert model responses into binary preference pairs using a consistent metric based on normalized benchmark performance. This approach entirely eliminates the need for instance-wise annotations, as it operates at the dataset level, transforming the entire instruction-response dataset into preference data without requiring instance-wise annotations from either models or humans. By leveraging aggregate performance metrics rather than per-instance comparisons, we maintain consistency while removing computational overhead and human cost.
2.4 Stage 3: Automatic Preference Integration with Alignment Processes
To integrate kinship-based mapped data into existing alignment tuning, the data requires a preference-setting process. Accordingly, we construct our kinship-based preference estimation on the principle that higher-performing models tend to produce more reliable outputs.
To quantify this, we calculate a normalized benchmark score that aggregates multiple benchmark performances into a single value. For a model , this score is computed as:
For a pair of kinship-aligned models , we obtain a binary preference label:
Our data maintains the Response Quality Gap through kinship-based model comparisons and can establish preferences via without additional annotation. This provides high-quality training signals, which can be leveraged through reinforcement learning or supervised fine-tuning.
In contrast to existing approaches (e.g., RLAIF Lee et al. ), KARMA’s focus on model kinship produces preference pairs that yield more nuanced insights. This shift in perspective—from relying on comparisons of models with vastly different capabilities to leveraging finely matched pairs—provides a more parsimonious and theoretically grounded method for constructing alignment-relevant preference data.
Baseline Ours RLAIF Total 0.31 0.31 0.29 0.29 MMLU 0.36 0.33 0.34 0.30 MMLU-Pro 0.15 0.15 0.15 0.14 ARC-Easy 0.40 0.40 0.36 0.41 ARC-Challenge 0.40 0.37 0.33 0.39 IFeval 0.28 0.30 0.29 0.23
3 Experiments
In this paper, we propose the KARMA framework. To validate the functionality of this framework and the characteristics of Alignment Tuning, we conducted experiments to address the following Research Questions (RQs):
RQ1: Does Model Kinship represent the models’ similarity?
RQ2: Does Model Kinship affect the binarization process?
RQ3: Does Model Kinship have a greater influence than model quality, as verified through benchmark evaluations?
RQ4: What benchmark represent the model capability: knowledge, instruction-following, and reasoning?
3.1 Alignment Dataset
For preference binarization, we utilized the Ultrafeedback dataset Cui et al. (2023), which comprises responses generated by a diverse set of language models. The dataset includes outputs from both commercial and open-source models, ensuring a broad representation of model capabilities. The commercial models incorporated in the dataset include GPT-4OpenAI (2023), GPT-3.5 TurboOuyang et al. (2022a), and BardWaisberg et al. (2024). Additionally, several models from the Llama family were included, such as Llama-2 (7B, 13B, and 70B)-chatTouvron et al. (2023), UltraLM-13BCui et al. (2024), WizardLM (7B, 13B, and 70B)Xu et al. (2023), Vicuna-33BZheng et al. (2024), and Alpaca-7BTaori et al. (2023). Beyond the Llama-based architectures, the dataset also features responses from other notable models, including Falcon-40B-instructAlmazrouei et al. (2023), MPT-30B-chatTeam (2023), StarChat-BetaTunstall et al. (2023), and Pythia-12BBiderman et al. (2023).
Although the Ultrafeedback dataset contains results from UltraLM-65BCui et al. (2024), its performance could not be accurately assessed. To maintain the reliability of our evaluation, we excluded these results from the final dataset composition.
3.2 Models
To assess the effectiveness of the newly binarized dataset constructed using the KARMA framework, we fine-tuned and evaluated multiple instruction-following models. The selected models for evaluation include Llama-3.1-(3B, 8B)-InstructDubey et al. (2024), representing Llama-based architectures. Additionally, we included Qwen2.5-(3B, 8B)-InstructYang et al. (2024), enabling a comparative analysis across different model families. This evaluation setup ensures a comprehensive assessment of the impact of kinship-aware preference mapping on alignment performance.
3.3 Training Algorithm for Alignment Tuning
We tested two different learning method for alignment tuning using the KARMA Framework:
Supervised Fine-Tuning (SFT)
Direct Preference Optimization (DPO) Rafailov et al. (2024)
The implementation details are provided in Appendix A
4 Experimental Results
4.1 RQ1: Dataset Reconstruction using KARMA
The proposed KARMA framework enables the application of a unified Total Ranking map across the entire dataset, facilitating a structured and consistent preference mapping process. Utilizing this reconstructed preference dataset, we conducted model training while ensuring that each critic’s binarized response was systematically incorporated. The effectiveness of this reconstructed dataset was assessed by evaluating the trained models on standardized benchmarks, including MMLU and IFeval. The detailed results of these evaluations are presented in Appendix C.
For evaluation, we employed prediction-based assessment methodologies across all benchmark tasks. Specifically, for ARC, MMLU, and MMLU-Pro, we adapted the MMLU-Pro evaluation framework, modifying only the multiple-choice options to align with our dataset. For IFeval, we leveraged its native evaluation framework to ensure consistency in assessment.
To analyze the relationships between models, we computed kinship scores by normalizing evaluation results across six benchmark tasks. Figure 4 presents a Principal Component Analysis (PCA)Maćkiewicz and Ratajczak (1993) visualization of these relationships, illustrating distinct clustering patterns among models. Generally, smaller-scale models tend to cluster in the first quadrant, models with closer kinship relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants. This distribution underscores the critical role of model training data, training algorithms, and scale in determining kinship relationships among models.
These findings highlight the effectiveness of the KARMA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
4.2 RQ2: Performance of KARMA Binarization
To assess the effectiveness of KARMA binarization, we examined whether model performance can serve as an indicator of data quality. KARMA performs binarization in two sequential steps: first, it defines preference pairs based on kinship between models, and second, it ranks these pairs according to their relative performance. Through this process, we aimed to determine whether response similarity—estimated through model similarity—plays a more significant role in preference modeling than simply prioritizing higher-performing models by consistently positioning them as dominant within each pair.
The results of this comparison are presented in Figure 5. Across all benchmark tasks, the kinship-based scoring metric, , demonstrates performance levels nearly equivalent to the existing instance-wise RLAIF method, with a minimal deviation of only 0.008. Notably, for MMLU-Pro and IFeval, KARMA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via kinship can yield competitive or superior alignment outcomes. The detailed results, including average scores for each methodology, are summarized in Appendix D.
4.3 RQ3: Model Size and Kinship
The impact of different kinship approaches varies significantly with model size, influencing how models learn from preference data. To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters). Our analysis reveals a clear pattern in how model size determines the optimal kinship strategy.
Smaller models (3B) exhibit superior performance when trained using the algorithm, which prioritizes learning from response quality within the binarized dataset. In contrast, larger models (7-8B) achieve better results with the algorithm, suggesting that response similarity becomes increasingly important as model size grows.
This trend indicates that model size fundamentally influences how different models leverage preference data. While smaller models benefit more from explicit learning based on absolute quality differences, larger models demonstrate greater sensitivity to the nuanced relationships between similar responses in the training data. This relationship is visualized in Figure 6, illustrating the distinct learning behaviors observed across different model scales.
4.4 RQ4: Model Capability and Kinship Patterns
Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics. Notably, model kinship is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
Analysis Framework
To systematically investigate these relationships, we analyzed model similarities across three key capability dimensions:
Knowledge-Based Tasks: Unlike reasoning capability, clustering in instruction-following tasks is more strongly aligned with model families rather than size. This indicates that training methodology and architectural choices exert a greater influence on instruction adherence.
Reasoning Capability: Models tend to cluster primarily based on parameter count, suggesting that model size plays a dominant role in shaping reasoning performance.
Instruction-Following (IF) Tasks: A hierarchical influence pattern emerges, where:
At the initial hierarchy, the model family is the primary determinant.
At the higher hierarchy, the model size becomes a stronger predictor of performance.
Figure 7 visualizes these relationships through dendrograms, illustrating the hierarchical clustering patterns that emerge across different capability dimensions.
5 Conclusion
This paper introduces KARMA, a kinship-based framework for generating high-quality preference data for language model alignment without requiring instance-wise annotation. KARMA achieves performance comparable to RLAIF while offering greater computational efficiency and surpassing baselines on MMLU-Pro and IFeval, demonstrating the effectiveness of kinship-aware preference mapping. Our findings show that model size significantly affects the optimal kinship strategy. Smaller models (3B) benefit from prioritizing response quality, while larger models (7-8B) perform better when leveraging response similarity. Additionally, kinship patterns vary across tasks, with reasoning clustering by model size, instruction-following by model family, and knowledge-based tasks exhibiting a hierarchical structure. These results highlight the need for size-adaptive preference learning strategies and demonstrate that high-quality preference data can be generated without costly manual annotation. Future work will extend KARMA to cross-family model comparisons and further explore the relationship between model kinship and human preference judgments.
6 Related Work and Background
6.1 Alignment Tuning
Alignment tuning has become a central focus in enhancing LLMs to meet user expectations and ethical standardsKumar et al. (2024). Various preference-based learning techniques, particularly reinforcement learning methods such as Proximal Policy Optimization (PPO)Schulman et al. (2017), Direct Preference Optimization (DPO)Rafailov et al. (2024), and Optimized Reinforcement Preference Optimization (ORPO)Hong et al. (2024), have been developed to facilitate this tuning process. These methods rely on preference datasets, which typically contain pairs of responses generated by models based on given instructions, with each pair ranked according to human or model-based evaluations. Prominent alignment datasets like UltrafeedbackCui et al. (2023), which gathers extensive human feedback, and HH-RLHFBai et al. (2022), which uses human-annotated preferences, provide foundational resources for alignment. To minimize the reliance on labor-intensive data curation, recent automated approaches have been introduced to filter or regenerate preference data based on specific criteria, promoting data consistency and scalability. However, such approaches often lack nuanced control over data quality, as they fail to consider the role of the origin model’s capabilities in shaping alignment effectiveness Shi et al. (2024).
6.2 Limitations in Existing Preference Data Approaches
Effective preference data construction requires a clear, rigorous set of criteria to ensure alignment quality across generated response pairs. Common criteria, including Reasoning, Truthfulness, and Instruction-Following, guide the selection of data that aligns with key ethical and functional standardsCui et al. (2023); Bai et al. (2022). High-performing models, capable of producing responses that meet these standards, are often evaluated using benchmarks like ARCClark et al. (2018), MMLUHendrycks et al. (2020), and the Instruction-Following evalZhou et al. (2023) suite, which assess a model’s factual accuracy, reasoning ability, and compliance with instructions. While these benchmarks provide a foundation for selecting responses that promote alignment goals, existing methods typically overlook the impact of variation in model capabilities on alignment consistency.
The KARMA Framework addresses this limitation by introducing a kinship-aware approach that emphasizes model compatibility in preference data selection. By curating preference pairs based on model kinship—aligning models with similar core competencies in knowledge, reasoning, and instruction-following—the KARMA Framework enhances alignment coherence and ensures stable, high-quality data. Further, its curriculum-based selection process sequences preference data from simpler to more complex distinctions, optimizing the alignment process for improved consistency and scalability.
Limitation
In this paper, we demonstrated that KARMA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO. However, due to resource and time constraints, we were unable to validate our approach across a broader range of alignment tuning techniques. Further evaluation is needed for methods.
Additionally, our evaluation primarily focused on fundamental abilities such as Reasoning, Knowledge, and Instruction-Following. However, we did not assess KARMA’s performance on other important values, including factuality and ethical standard. Future work should incorporate evaluations reflecting these aspects.
Appendix A Implementation Details for Model Training
The model was trained for a single epoch using bfloat16 (bf16) quantization, which optimizes memory efficiency while preserving numerical precision. The training configuration incorporated the following hyperparameters: a per-device batch size of 6, gradient accumulation steps set to 4, and a learning rate of , with 500 warm-up steps to facilitate stable convergence.
Training was conducted on an L40 4-GPU setup, leveraging an optimized deep learning framework to enhance computational efficiency. The training pipeline focused on performance optimization through checkpointing and logging, without intermediate model evaluation during training.
Appendix B Prompt template for Evaluation Benchmarks
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 8. For the start and end tokens, the template recommended in the paper was used.
Appendix C Evaluation Origin Model Capabilities for Kinship
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 3. The actual calculation example for and can be found in Figure 9.
Due to the unavailability of Bard, its performance metrics have been substituted with those of Gemini-1.5-Flash.
Model IFeval MMLU-STEM MMLU-pro Hellaswag ARC-easy ARC-Challenge Average Score GPT-4 85.37 86.40 0.64 95.30 96.63 96.40 0.99 GPT-3.5-turbo 72.54 70.00 0.38 85.00 92.80 83.02 0.77 bard 89.33 71.80 0.59 84.70 84.43 77.13 0.86 Llama-2-7b-chat 25.19 53.10 0.20 67.50 72.14 54.61 0.41 Llama-2-13b-chat 24.82 57.80 0.25 71.20 72.05 58.02 0.45 Llama-2-70b-chat 24.07 68.90 0.38 78.10 82.20 67.66 0.58 UltraLM-13b 54.92 49.58 0.19 54.00 57.58 48.04 0.40 WizardLM-7b 45.83 42.50 0.18 77.70 39.48 34.90 0.34 WizardLM-13b 33.92 52.30 0.17 81.00 72.94 55.38 0.45 WizardLM-70b 49.51 52.70 0.27 83.30 80.68 71.93 0.58 Vicuna-33b 52.76 64.00 0.23 75.00 81.57 64.51 0.57 Alpaca-7b 30.58 37.92 0.15 23.60 43.31 33.79 0.16 Falcon-40b-instruct 24.54 67.50 0.14 80.00 76.60 56.70 0.47 MPT-30b-chat 30.70 50.40 0.20 24.53 87.12 70.73 0.38 Starchat 28.30 40.12 0.12 25.40 16.96 9.07 0.05 Pythia-12b 24.71 27.00 0.12 25.60 24.49 31.80 0.07
Appendix D Total Performance
Table 4 provides an aggregated view of the overall performance of each model across all benchmark datasets. The total performance scores were computed by averaging the normalized scores across the selected evaluation metrics, offering a holistic comparison of model capabilities.
| Benchmark | Model | Baseline | ||||
| MMLU-STEM | Llama-3.1-3B | SFT | 0.2331 | 0.2679 | 0.2434 | 0.2200 |
| DPO | 0.3130 | 0.4868 | 0.2528 | 0.3940 | ||
| Llama-3.1-8B | SFT | 0.2800 | 0.2460 | 0.2660 | 0.2240 | |
| DPO | 0.4902 | 0.4349 | 0.4867 | 0.2880 | ||
| Qwen2.5-3B | SFT | 0.2760 | 0.2200 | 0.2460 | 0.2500 | |
| DPO | 0.4706 | 0.4212 | 0.4400 | 0.4580 | ||
| Qwen2.5-7B | SFT | 0.2800 | 0.2280 | 0.2620 | 0.2780 | |
| DPO | 0.5120 | 0.3620 | 0.4860 | 0.2800 | ||
| MMLU-pro | Llama-3.1-3B | SFT | 0.1189 | 0.1230 | 0.1045 | 0.1332 |
| DPO | 0.1455 | 0.2193 | 0.1270 | 0.1516 | ||
| Llama-3.1-8B | SFT | 0.1004 | 0.1148 | 0.1045 | 0.1311 | |
| DPO | 0.1168 | 0.1025 | 0.1168 | 0.1107 | ||
| Qwen2.5-3B | SFT | 0.1025 | 0.0840 | 0.1230 | 0.0820 | |
| DPO | 0.2275 | 0.2254 | 0.2029 | 0.2111 | ||
| Qwen2.5-7B | SFT | 0.0861 | 0.1762 | 0.2275 | 0.1352 | |
| DPO | 0.2377 | 0.1721 | 0.2254 | 0.1700 | ||
| IFeval | Llama-3.1-3B | SFT | 0.2494 | 0.2410 | 0.2490 | 0.2206 |
| DPO | 0.3765 | 0.4940 | 0.4796 | 0.3033 | ||
| Llama-3.1-8B | SFT | 0.2494 | 0.4210 | 0.2470 | 0.2218 | |
| DPO | 0.2421 | 0.1882 | 0.2292 | 0.2494 | ||
| Qwen2.5-3B | SFT | 0.2190 | 0.2134 | 0.2122 | 0.2050 | |
| DPO | 0.3633 | 0.3058 | 0.3094 | 0.1715 | ||
| Qwen2.5-7B | SFT | 0.2290 | 0.2134 | 0.2122 | 0.2083 | |
| DPO | 0.3321 | 0.3177 | 0.3657 | 0.2407 | ||
| ARC-easy | Llama-3.1-3B | SFT | 0.2660 | 0.2460 | 0.2000 | 0.2340 |
| DPO | 0.6111 | 0.5547 | 0.2618 | 0.6679 | ||
| Llama-3.1-8B | SFT | 0.2180 | 0.2330 | 0.2400 | 0.2360 | |
| DPO | 0.2176 | 0.1540 | 0.4720 | 0.0680 | ||
| Qwen2.5-3B | SFT | 0.2410 | 0.2560 | 0.2480 | 0.2560 | |
| DPO | 0.8497 | 0.8754 | 0.5295 | 0.8157 | ||
| Qwen2.5-7B | SFT | 0.2550 | 0.2260 | 0.2280 | 0.2900 | |
| DPO | 0.5700 | 0.6359 | 0.6738 | 0.7134 | ||
| ARC-challenge | Llama-3.1-3B | SFT | 0.2556 | 0.2492 | 0.2266 | 0.2019 |
| DPO | 0.5122 | 0.4551 | 0.2466 | 0.5712 | ||
| Llama-3.1-8B | SFT | 0.2320 | 0.2297 | 0.2761 | 0.2268 | |
| DPO | 0.5463 | 0.1763 | 0.3596 | 0.2946 | ||
| Qwen2.5-3B | SFT | 0.2645 | 0.2483 | 0.2227 | 0.2343 | |
| DPO | 0.7398 | 0.8241 | 0.4470 | 0.7078 | ||
| Qwen2.5-7B | SFT | 0.2343 | 0.2552 | 0.2483 | 0.2390 | |
| DPO | 0.4432 | 0.5358 | 0.5847 | 0.3870 |

### v2

@@
-Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA) for Automatic Alignment in Large Language Models
+ZEBRA: Leveraging Model-Behavioral Knowledge for Zero-Annotation Preference Dataset Construction
 Abstract
-Recent advancements in Large Language Model (LLM) alignment have sought to mitigate the cost of human annotations by leveraging pretrained models to generate preference data.
-However, existing methods often compare responses from models with substantially different capabilities, yielding superficial distinctions that fail to provide meaningful guidance on what constitutes a superior response.
-To address this limitation, we propose Kinship-Aware pReference MApping (KARMA), a novel framework that systematically pairs responses from models with comparable competencies.
-By constraining preference comparisons to outputs of similar complexity and quality, KARMA enhances the informativeness of preference data and improves the granularity of alignment signals.
-Empirical evaluations demonstrate that our kinship-aware approach leads to more consistent and interpretable alignment outcomes, ultimately facilitating a more principled and reliable pathway for aligning LLM behavior with human preferences.
-Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA)
-for Automatic Alignment in Large Language Models
-Jeesu Jung1, Chanjun Park2, Sangkeun Jung1††thanks: Corresponding author 1Chungnam National University, 2Korea University jisu.jung5@gmail.com, bcj1210@korea.ac.kr, hugmanskj@gmail.com
+Recent efforts in LLM alignment have focused on constructing large-scale preference datasets via human or Artificial Intelligence(AI) annotators.
+However, such approaches rely on instance-wise supervision, incurring substantial annotation cost and limited interpretability.
+In this paper, we propose ZEBRA—a model behavior-wise zero-annotation framework that constructs preference data by leveraging model behavior knowledge derived from benchmark performances.
+ZEBRA binarizes response pairs by evaluating the quality and similarity of their origin models, entirely bypassing instance-level annotation.
+This allows scalable, controllable, and cost-effective alignment data generation.
+Empirical results show that ZEBRA achieves alignment performance comparable to instance-supervised methods, despite requiring no manual or model-based labeling.
+ZEBRA: Leveraging Model-Behavioral Knowledge
+for Zero-Annotation Preference Dataset Construction
+Jeesu Jung1, Chanjun Park2⋆, Sangkeun Jung1††thanks: Equal contribution.
+Corresponding author 1Chungnam National University, 2Korea University jisu.jung5@gmail.com, bcj1210@korea.ac.kr, hugmanskj@gmail.com
 1 Introduction
-Aligning Large Language Models (LLMs) with human preferences is a fundamental challenge in artificial intelligence, where the effectiveness of alignment hinges on the quality and specificity of the preference data used to guide model outputs Shen et al.
-.
-Early alignment methodologies primarily relied on human-annotated datasets with binary preference labels, wherein explicit human judgments directed models toward generating more desirable responses Ouyang et al.
-(2022a).
-While these methods have played a critical role in improving LLM behavior, their dependence on extensive human labor imposes significant scalability constraints.
-To address this limitation, recent research has explored automated preference data generation techniques, such as Reinforcement Learning from AI Feedback (RLAIF) Lee et al.
-.
-These methods leverage pretrained models to automatically generate preference pairs, significantly reducing the need for human annotation.
-However, the quality of the resulting preference data remains a major concern.
-If responses within a preference pair exhibit substantial disparities in quality, complexity, or style, the preference signal may become trivial, capturing only superficial contrasts rather than informative distinctions that refine model behavior.
-We define this issue as the Response Quality Gap, which refers to the disparity in coherence, complexity, and reasoning quality between candidate responses.
-A large Response Quality Gap weakens preference signals, leading to misaligned or less informative training objectives.
-This problem is particularly pronounced in large-scale, automatically generated datasets, where manually curating response pairs to minimize the Response Quality Gap is infeasible due to the massive volume of data.
-A promising solution lies in considering the origin of the responses being compared—specifically, the LLM that generated them.
-Since a model’s reasoning ability, commonsense knowledge, and instruction-following capabilities are inherently shaped by its training data and architecture, the quality of its responses is largely determined by these characteristics.
-Consequently, comparing responses from models with vastly different capabilities introduces inconsistencies in preference data, making alignment objectives less reliable.
-To formalize this idea, we introduce the notion of Kinship among LLMs, which quantifies the similarity in reasoning ability, commonsense understanding, and instruction adherence among models.
-Measured using standard benchmarks, kinship provides a principled way to assess whether two models generate responses of comparable complexity and quality.
-Building on this insight, we propose the Kinship-Aware pReference MApping (KARMA) framework, which systematically pairs responses based on model kinship.
-Unlike conventional approaches that indiscriminately compare outputs from models of varying capabilities, KARMA ensures that preference pairs are drawn from models with similar competencies, thereby narrowing the Response Quality Gap.
-This results in preference data that captures more subtle and meaningful distinctions, leading to higher-quality alignment signals.
-The effectiveness of KARMA is illustrated in Figure 1, which provides an overview of the framework, and Figure 2, which details its response evaluation process.
-By fostering more structured and granular comparisons, KARMA improves the coherence and interpretability of preference data, shifting alignment objectives from filtering out trivial differences to refining preference signals through comparability and consistency.
-Our empirical results demonstrate that this kinship-aware approach significantly outperforms prior methods such as RLAIF, offering a more robust and scalable solution for aligning LLM behavior with human preferences.
-2 Kinship-Aware pReference MApping (KARMA) Framework
-Recent approaches to LLM alignment have reduced the reliance on human annotation through automated preference data generation.
-However, these methods often neglect the foundational influence of the Origin Model.
-For example, a model trained using Llama’s methodology is going to produce outputs that fall within the distribution of Llama’s training data.
-Just as a son inherits traits from his father, a model’s quality directly influences the quality of its outputs.
-when preference pairs are derived from models with markedly disparate capabilities, the resulting comparisons yield only trivial distinctions.
-To address this, the proposed Kinship-Aware pReference MApping (KARMA) framework evaluates kinship based on the performance of origin models on selected benchmark metrics.
-With kinship as a guiding principle for structuring binarized data, the framework systematically generates preference data for alignment tuning.
-2.1 Definition of Kinship
-In the context of binarized alignment data mapping, the similarity between the responses serves as a fundamental determinant of mapping efficacy.
-When data pairs are closely aligned, they provide more nuanced and meaningful guidance signals for alignment tuning.
-Building upon this principle of response similarity, KARMA introduces the concept of kinship among models.
-The kinship is measured by evaluating the capabilities of origin models.
-We measure the benchmarks as the origin model capability.
-We define a set of candidate origin models, , each evaluated according to multiple dimensions of competency.
-Let be a set of benchmark tasks.
-Each benchmark provides an evaluation metric that quantifies the performance of model .
-This evaluation may include measures of factual correctness, reasoning, and adherence to instructions.
-For each pair of models , we define a kinship score , which reflects the similarity of their capabilities:
-where is a similarity function.
-Higher values of indicate closer competency alignment between and .
-The exact functional form of can vary depending on the chosen benchmarks and may incorporate multiple factors (denoted collectively as , , and ) to minimize bias and ensure comprehensive evaluation.
-Descriptive Multiple-Choice Instruction Following IF evalZhou et al.
-(2023) Knowledge MMLUHendrycks et al.
-(2020) MMLU-proWang et al.
-(2024) Reasoning ARC-EasyClark et al.
-(2018) ARC-ChallengeClark et al.
-(2018) HellaswagZellers et al.
-(2019)
-2.2 Stage 1: Benchmark Evaluation for Model Competency
-To identify which models are suitable candidates for kinship-based pairing, we employ a series of benchmarks that meet three criteria: (1) they are widely validated in prior research, (2) they cover both general knowledge and domain-specific tasks, and (3) they provide well-defined quantitative metrics.
-A summary of selected benchmarks is provided in Table 1.
-Evaluating models on these benchmarks yields performance scores that serve as the basis for computing .
-The degree of kinship among models varies with the choice of benchmarks.
-To avoid relying solely on log-probabilities or overly simplistic metrics, we adopt prompt-based evaluations that require each model to produce explicit responses.
-Appendix C presents the benchmark performance of all models used for kinship calculation, along with example calculations.
-The prompt formats are adapted from templates in MMLU-pro Wang et al.
-(2024) and FLAN Longpre et al.
-(2023), ensuring compatibility with each model’s input structure and recommended usage patterns.
-The prompt template can be found in Appendix B.
-2.3 Stage 2: Kinship-Based Response Mapping
-Once we identify a subset of models whose pairwise kinship scores exceed a predetermined threshold (e.g., ), we set at 0.1, as cosine similarity values below this threshold indicate a lack of meaningful similarity between models.
-we focus on collecting responses from these closely matched models.
-Kinship serves as an indicator of model’s ability and similarity.
-While various criteria can be applied to measure this, we utilize the models’ capabilities and performance similarities that can be estimated through benchmarks.
-We have defined kinship across three distinct cases:
-: Mapping the two origin models with the highest average benchmark scores.
-: Mapping the two origin models with the highest cosine similarity between their benchmark scores.
-: Mapping the model with the highest benchmark score to the most similar model based on cosine similarity.
-Figure 3 shows the each kinship definition and calculation process.
-To generate preference data from these kinship relationships, we convert model responses into binary preference pairs using a consistent metric based on normalized benchmark performance.
-This approach entirely eliminates the need for instance-wise annotations, as it operates at the dataset level, transforming the entire instruction-response dataset into preference data without requiring instance-wise annotations from either models or humans.
-By leveraging aggregate performance metrics rather than per-instance comparisons, we maintain consistency while removing computational overhead and human cost.
-2.4 Stage 3: Automatic Preference Integration with Alignment Processes
-To integrate kinship-based mapped data into existing alignment tuning, the data requires a preference-setting process.
-Accordingly, we construct our kinship-based preference estimation on the principle that higher-performing models tend to produce more reliable outputs.
-To quantify this, we calculate a normalized benchmark score that aggregates multiple benchmark performances into a single value. For a model , this score is computed as:
-For a pair of kinship-aligned models , we obtain a binary preference label:
-Our data maintains the Response Quality Gap through kinship-based model comparisons and can establish preferences via without additional annotation.
-This provides high-quality training signals, which can be leveraged through reinforcement learning or supervised fine-tuning.
-In contrast to existing approaches (e.g., RLAIF Lee et al.
-), KARMA’s focus on model kinship produces preference pairs that yield more nuanced insights.
-This shift in perspective—from relying on comparisons of models with vastly different capabilities to leveraging finely matched pairs—provides a more parsimonious and theoretically grounded method for constructing alignment-relevant preference data.
-Baseline Ours RLAIF Total 0.31 0.31 0.29 0.29 MMLU 0.36 0.33 0.34 0.30 MMLU-Pro 0.15 0.15 0.15 0.14 ARC-Easy 0.40 0.40 0.36 0.41 ARC-Challenge 0.40 0.37 0.33 0.39 IFeval 0.28 0.30 0.29 0.23
-3 Experiments
-In this paper, we propose the KARMA framework.
+Aligning large language models (LLMs) with human preferences is an essential step toward making them both useful and safe.
+A common way to achieve this is through instance-wise labeling, where pairs of model responses are compared one by one to see which is better.
+Well-known methods like Reinforcement Learning from Human Feedback (RLHFOuyang et al.
+(2022a)) and Artificial Inteligence(AI)-based labeling(RLAIFLee et al.
+) often use this strategy.
+However, instance-wise labeling faces two major challenges.
+First, it is very costly, whether it involves human annotators or additional computational resources for LLM-based labelingZhang et al.
+(2024); Zheng et al.
+(2023).
+Second, it lacks a global view of the model’s behaviorJi et al.
+(2023).
+Since each response pair is judged in isolation, it is difficult to consider broader factors.
+For example, whether fluency should outweigh factual accuracy.
+Or whether a model’s outputs are consistently aligned with certain policiesWang et al.
+(2025).
+This can lead to labeling noise, mistakes, and limited interpretability.
+To address these limitations, we propose a new preference binarization approach called ZEro-annotation Behavior-based Response Alignment (ZEBRA) (Figure 1).
+The main idea is: (1) extract each model’s behavioral patterns from its past performance trajectories, (2) measure and compare these behaviors in terms of model strength or similarity, and (3) assign preferences at the model level rather than for each individual response pair.
+We implement this idea through three key components.
+First, we define Model Behavior Knowledge (MBK) for LLMs (discuss in Section 4.1).
+Second, we propose a way to quantify and collect MBK from objective data sources such as benchmark performance.
+Third, we introduce three strategies—based on superiority, similarity, and a hybrid—to construct a binarization dataset in a zero-annotation, cost-free manner.
+A major advantage of our approach is that it creates response pairs based on model superiority without any additional human or LLM labeling.
+By classifying models with higher benchmark scores as “positive” and those with lower scores as “negative,” we can systematically label preferences throughout the dataset.
+This significantly reduces the cost of annotation and, because MBK visualizes each model’s behavior pattern, increases the interpretability of the preference decisions.
+Through extensive experiments, we show that ZEBRA achieves performance comparable to existing instance-wise labeling methods in the UltrafeedbackCui et al.
+(2023) dataset—without any extra labeling cost.
+In summary, our contributions are as follows:
+We introduce ZEBRA, a zero-annotation alignment framework that determines preferences from quantified model-level behavior, bypassing instance-level supervision.
+We demonstrate that the Model Behavior Knowledge (MBK) from benchmark performance offers alignment signals comparable to instance-wise labeling.
+We empirically show that ZEBRA matches the performance of established methods such as RLHF and RLAIF, yet requires no additional labeling cost.
+2 Preliminaries
+2.1 Instance-Level Preference Construction
+Most existing preference learning frameworks for LLM alignment—such as Reinforcement Learning from Human Feedback(RLHF, Ouyang et al.
+(2022a)) and AI-generated preference methods (RLAIF, Lee et al.
+)—rely on instance-level pairwise supervision.
+Given an instruction , multiple candidate responses are scored or ranked by either human annotators or automated scoring models.
+This generates preference tuples , where is preferred over under some evaluation criteria (e.g., helpfulness, truthfulness, coherence).
+The preference construction process typically involves:
+Generating multiple responses per instruction using different models or decoding strategies.
+Computing preference labels via either human judgment or model-based scoring (e.g., GPT-4OpenAI (2023)).
+Aggregating these labels into a pairwise dataset for alignment tuning.
+2.2 Challenges of Instance-Level Supervision
+While instance-level supervision has proven effective in aligning LLMs, it remains costly, noisy, and difficult to scale.
+Despite its popularity, instance-level supervision suffers from three core limitations:
+1.
+Costly evaluation: Annotating or scoring each response pair requires substantial human or computational effort.
+2.
+Preference triviality: When candidate responses differ significantly in quality, the preference label becomes trivial, contributing little to alignment learning.
+3.
+Instruction-level variance: Difficulty and ambiguity in the instruction can introduce noise into the preference signal, especially for automated labelers.
+These limitations underscore the need for alternative approaches that construct preference signals without relying on per-instance scoring.
+2.3 Motivation for Model Behavior-Level Preference
+We propose that instead of relying on per-instance evaluation, one can leverage the intrinsic capabilities of the response-generating models themselves.
+As models exhibit differences in core competencies measurable via standardized benchmarks, we hypothesize this can replace instance-level annotations.
+3 ZEro-annotation Behavior-based Response Alignment Framework
+3.1 Instance-level Annotation vs.
+Model behavior-level Annotation
+Traditional instance-level binarization relies heavily on detailed, provided by human or AI annotatorsZhang et al.
+(2024); Sharma et al.
+(2024).
+Each pairwise comparison demands significant effort and resources to maintain consistency and interpretability.
+Such an approach often results in annotation noise, limited scalability, and considerable expense.
+In contrast, our proposed ZEBRA framework leverages intrinsic behavioral knowledge derived from model performance across various benchmarks, ZEBRA systematically matches and pairs responses.
+Responses from models with proven higher competencies form positive labels, while those from models with lower competencies become negative labels.
+This innovative model-level approach drastically reduces annotation costs, minimizes noise, and enhances scalability.
+Additionally, the behavior similarity-based matching provides explicit control over the difficulty and nuance of preference comparisons, leading to clearer and more meaningful alignment outcomes.
+3.2 Model Behavior Knowledge
+While most preference-learning pipelines focus on differences between individual responses, our approach highlights Model Behavior Knowledge (MBK)—the comprehensive record of each model’s past behaviors and capabilities.
+We define MBK using two sets of metrics:
+Superiority: How much better (or worse) a model is compared to others, based on overall or task-specific proficiency.
+Similarity: How likely a model is to behave similarly to other models.
+These metrics provide a principled basis for quantifying both a model’s general strength and its behavioral proximity to its peers.
+Within a preference-learning pipeline, superiority functions as a global preference signal: when one model consistently outperforms another across standardized benchmarks, its responses are considered preferable overall.
+Conversely, behavioral similarity facilitates the systematic construction of challenging comparison sets.
+When two models are behaviorally similar—for example, they exhibit comparable reasoning performance—their responses become difficult to distinguish.
+Training on such hard-to-distinguish pairs guides the preference learner to focus on subtle qualitative differences, resulting in more nuanced and robust alignment.
+3.3 Model Behavior Evaluation using Benchmark Performances
+To capture MBK in a practical, objective way, we rely on external benchmark performance data for each LLM.
+Many models are already evaluated across diverse, standardized tasks (e.g., reasoning, factual accuracy, instruction-following).
+We aggregate these benchmark scores to form each model’s MBK profile.
+Benchmark performance offers several advantages for extracting MBK:
+It provides reliable metrics for the core competencies of large language models.
+It enables straightforward comparison and aggregation across multiple models.
+Because benchmark scores are published and fixed at release, they impose no constraints on subsequent data or model expansion.
+The list of benchmarks used in our analysis is provided in Figure 3.
+Figure 4 illustrates how the behavioral similarities inferred from benchmark performance can reflect actual model similarities.
+For example, the LLaMA-2-13b model exhibits a pattern closely resembling that of its 7b counterpart.
+In contrast, WizardLM-7b demonstrates a markedly different behavioral trajectory.
+We define a model’s ability vector across benchmark tasks:
+where is the normalized score of model on benchmark , reflecting its relative capability in a specific behavioral dimension (e.g., knowledge, reasoning, instruction-following).
+These standardized behavior vectors serve as the foundation for both quality-based and similarity-based anchoring, enabling zero-annotation binarization of preference pairs without per-instance supervision.
+3.4 Benchmark-based Model Behavior Quantification
+To effectively binarize preference data, it is essential to quantify model behavior from two complementary perspectives: behavior quality and behavior similarity.
+ZEBRA leverages benchmark-derived measures of these aspects to systematically pair positive responses with suitable negative counterparts.
+By quantifying both the absolute competency of individual models and the relative similarity between models, ZEBRA ensures that each positive-negative pair captures meaningful contrasts in model capabilities, thus maximizing alignment informativeness.
+Model Behavior Superiority (MB-SUP) quantifies the overall behavioral competency of model as the aggregate of its normalized benchmark scores:
+This scalar value serves as the ranking basis for constructing Behavioral Superiority Anchors.
+Model Behavior Similarity (MB-SIM) between models and is defined as the similarity between their behavior vectors:
+Higher values indicate stronger alignment in general-purpose capabilities, and MB-SIM serves as the criterion for selecting comparable model pairs in Behavioral Similarity Anchoring.
+3.5 Strategy of Preference Binarization
+ZEBRA introduces multiple strategies to systematically convert benchmark-derived model behaviors into binary preference pairs.
+Based on how MB-SUP and MB-SIM are utilized, our strategies can be clearly categorized as follows on Figure 5.
+We detail each strategy below:
+Strategy 1: Superiority-first Anchoring (SUP)
+In this strategy, we explicitly select the top-two models based on their MB-SUP scores: the highest-scoring model (top-1) and the second-highest-scoring model (top-2).
+Responses from the top-1 model serve as positive anchors, while responses from the top-2 model become negative counterparts.
+This approach emphasizes explicit quality distinctions, clearly defining superior responses and ensuring meaningful, informative alignment contrasts.
+Strategy 2: Similarity-first Anchoring (SIM)
+Responses from models sharing similar behavioral patterns (high MB-SIM) are paired first.
+Within these pairs, the response from the model with higher MB-SUP is selected as the anchor (positive response).
+This strategy emphasizes behavioral similarity, enhancing nuanced alignment comparisons.
+Strategy 3: Hybrid Anchoring (SUP+SIM)
+Model pairs are selected by simultaneously considering both MB-SUP and MB-SIM criteria.
+This balanced approach ensures each response pair reflects meaningful contrasts in model quality, while maintaining behavioral similarity for refined granularity.
+These strategies enable ZEBRA to flexibly and effectively tailor preference construction according to the desired granularity, alignment objectives, and computational resources.
+Figure 5 represent the example of each strategy.
+3.6 Zero-Annotation Preference Construction
+Given a set of instructions and a pool of response-generating models , the construction proceeds as follows:
+1.
+Model Behavior Evaluation: Each model is benchmarked across tasks to obtain model’s ability vector .
+2.
+Pair Selection: Using a chosen strategy, a set of model pairs is selected where .
+we set at 0.1, as cosine similarity values below this threshold indicate a lack of meaningful similarity between models.
+3.
+Response Mapping: For each instruction , the corresponding responses from are retrieved.
+4.
+Preference Assignment: A binary label is assigned via:
+This pipeline constructs a binarized preference dataset at scale without any manual or per-instance scoring. Figure 2 shows the total process of the pipeline.
+4 Experimental Setup
+In this paper, we propose the ZEBRA framework.
 To validate the functionality of this framework and the characteristics of Alignment Tuning, we conducted experiments to address the following Research Questions (RQs):
-RQ1: Does Model Kinship represent the models’ similarity?
-RQ2: Does Model Kinship affect the binarization process?
-RQ3: Does Model Kinship have a greater influence than model quality, as verified through benchmark evaluations?
-RQ4: What benchmark represent the model capability: knowledge, instruction-following, and reasoning?
+RQ1: Does MB-SIM represent the models’ similarity?
+RQ2: Does strategies affect the binarization process?
+RQ3: Does ZEBRA reduce annotation cost and computational overhead compared to instance-level methods?
 4.1 Alignment Dataset
 For preference binarization, we utilized the UltraFeedback dataset Cui et al.
-(2023), which comprises responses generated by a diverse set of language models.
-The dataset includes outputs from both commercial and open-source models, ensuring a broad representation of model capabilities.
-The commercial models incorporated in the dataset include GPT-4OpenAI (2023), GPT-3.5 TurboOuyang et al.
-(2022a), and BardWaisberg et al.
+(2023), which includes diverse responses generated by both commercial and open-source language models.
+It provides multiple model responses, making it suitable for constructing a strong RLAIF baseline, for which we adopted the original score aggregation method.
+In contrast, ZEBRA uses only the response pool without referencing any scores, ensuring a clean comparison focused on the binarization strategy.
+Model coverage.
+UltraFeedback contains outputs from 17 LLMs: GPT-4OpenAI (2023), GPT-3.5 TurboOuyang et al.
+(2022b), and BardWaisberg et al.
 (2024).
@@
 (2024), its performance could not be accurately assessed.
-To maintain the reliability of our evaluation, we excluded these results from the final dataset composition.
+To maintain the reliability of our evaluation, we excluded these results from the dataset composition.
 4.2 Models
-To assess the effectiveness of the newly binarized dataset constructed using the KARMA framework, we fine-tuned and evaluated multiple instruction-following models.
-The selected models for evaluation include Llama-3.1-(3B, 8B)-InstructDubey et al.
-(2024), representing Llama-based architectures.
-Additionally, we included Qwen2.5-(3B, 8B)-InstructYang et al.
-(2024), enabling a comparative analysis across different model families.
-This evaluation setup ensures a comprehensive assessment of the impact of kinship-aware preference mapping on alignment performance.
 4.3 Training Algorithm for Alignment Tuning
-We tested two different learning method for alignment tuning using the KARMA Framework:
+We tested two different learning method for alignment tuning using the ZEBRA Framework:
 Supervised Fine-Tuning (SFT)
 Direct Preference Optimization (DPO) Rafailov et al. (2024)
-The implementation details are provided in Appendix A
+The implementation details are provided in Appendix A.
+The full source code, benchmark results, and preference binarization scripts will be released after publication via our project repository under the MIT License.
+The snapshot corresponding to this paper is version-tagged.
 5 Experimental Results
-4.1 RQ1: Dataset Reconstruction using KARMA
-The proposed KARMA framework enables the application of a unified Total Ranking map across the entire dataset, facilitating a structured and consistent preference mapping process.
+Knowledge Reasoning Instruction-Following Category Strategy Average MMLU MMLU-Pro ARC-Easy ARC-Challenge IFeval Baseline RLAIF 0.31 0.36 0.15 0.40 0.40 0.28 ZEBRA(Ours) SUP 0.31(-0.00) 0.33 0.15 0.40 0.37 0.30 SIM 0.29(-0.02) 0.34 0.15 0.36 0.33 0.29 SUP+SIM 0.29(-0.02) 0.30 0.14 0.41 0.39 0.23
+Comparison Mean Δ SUP vs BL -0.3541 0.7282 -0.0066 0.7345 SIM vs BL -0.4374 0.6680 -0.0047 0.7585 SUP+SIM vs BL -1.4472 0.1684 -0.0296 0.2011
+Category Method Pairs / Units Unit Cost (USD) Total Cost (USD) Instance-wise RLAIF UltraFeedbackCui et al.
+(2023) 64,000 0.252 16,128 Safer-InstructShi et al.
+(2024) 10,254 0.063 646 OpenHermesPreferencesHuang et al.
+(2024) 989,000 0.126 124,614 Model behavior-wise ZEBRA (ours) 64,000 0 0
+5.1 RQ1: Dataset Reconstruction using ZEBRA
+The proposed ZEBRA framework enables the application of a unified Total Ranking map across the entire dataset, facilitating a structured and consistent preference mapping process.
 Utilizing this reconstructed preference dataset, we conducted model training while ensuring that each critic’s binarized response was systematically incorporated.
-The effectiveness of this reconstructed dataset was assessed by evaluating the trained models on standardized benchmarks, including MMLU and IFeval.
+The effectiveness of this reconstructed dataset was assessed by evaluating the trained models on standardized benchmarks.
 The detailed results of these evaluations are presented in Appendix C.
@@
 For IFeval, we leveraged its native evaluation framework to ensure consistency in assessment.
-To analyze the relationships between models, we computed kinship scores by normalizing evaluation results across six benchmark tasks.
+To analyze the relationships between models, we computed SUP and SIM by normalizing evaluation results across six benchmark tasks.
 Figure 6 presents a Principal Component Analysis (PCA)Maćkiewicz and Ratajczak (1993) visualization of these relationships, illustrating distinct clustering patterns among models.
-Generally, smaller-scale models tend to cluster in the first quadrant, models with closer kinship relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants.
-This distribution underscores the critical role of model training data, training algorithms, and scale in determining kinship relationships among models.
-These findings highlight the effectiveness of the KARMA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
-4.2 RQ2: Performance of KARMA Binarization
-To assess the effectiveness of KARMA binarization, we examined whether model performance can serve as an indicator of data quality.
-KARMA performs binarization in two sequential steps: first, it defines preference pairs based on kinship between models, and second, it ranks these pairs according to their relative performance.
-Through this process, we aimed to determine whether response similarity—estimated through model similarity—plays a more significant role in preference modeling than simply prioritizing higher-performing models by consistently positioning them as dominant within each pair.
-The results of this comparison are presented in Figure 5.
-Across all benchmark tasks, the kinship-based scoring metric, , demonstrates performance levels nearly equivalent to the existing instance-wise RLAIF method, with a minimal deviation of only 0.008.
-Notably, for MMLU-Pro and IFeval, KARMA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via kinship can yield competitive or superior alignment outcomes.
-The detailed results, including average scores for each methodology, are summarized in Appendix D.
-4.3 RQ3: Model Size and Kinship
-The impact of different kinship approaches varies significantly with model size, influencing how models learn from preference data.
-To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters).
-Our analysis reveals a clear pattern in how model size determines the optimal kinship strategy.
-Smaller models (3B) exhibit superior performance when trained using the algorithm, which prioritizes learning from response quality within the binarized dataset.
-In contrast, larger models (7-8B) achieve better results with the algorithm, suggesting that response similarity becomes increasingly important as model size grows.
-This trend indicates that model size fundamentally influences how different models leverage preference data.
-While smaller models benefit more from explicit learning based on absolute quality differences, larger models demonstrate greater sensitivity to the nuanced relationships between similar responses in the training data.
-This relationship is visualized in Figure 6, illustrating the distinct learning behaviors observed across different model scales.
-4.4 RQ4: Model Capability and Kinship Patterns
-Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics.
-Notably, model kinship is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
-Analysis Framework
-To systematically investigate these relationships, we analyzed model similarities across three key capability dimensions:
-Knowledge-Based Tasks: Unlike reasoning capability, clustering in instruction-following tasks is more strongly aligned with model families rather than size.
-This indicates that training methodology and architectural choices exert a greater influence on instruction adherence.
-Reasoning Capability: Models tend to cluster primarily based on parameter count, suggesting that model size plays a dominant role in shaping reasoning performance.
-Instruction-Following (IF) Tasks: A hierarchical influence pattern emerges, where:
-At the initial hierarchy, the model family is the primary determinant.
-At the higher hierarchy, the model size becomes a stronger predictor of performance.
-Figure 7 visualizes these relationships through dendrograms, illustrating the hierarchical clustering patterns that emerge across different capability dimensions.
+Generally, smaller-scale models tend to cluster in the first quadrant, models with closer Model Behavior relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants.
+This distribution underscores the critical role of model training data, training algorithms, and scale in determining Model Behavior relationships among models.
+We further verify that anchor pairs selected by SIM indeed yield semantically closer responses.
+Using a TF–IDFRamos (2003), we obtain an average response-pair similarity of 0.4623 for the most similar model pair () versus 0.4129 for the least similar model pair ().
+This 12% relative gap confirms that high-MB-SIM anchoring indeed surfaces finer-grained yet coherent preference signals.
+These findings highlight the effectiveness of the ZEBRA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
+5.2 RQ2: Performance of ZEBRA Binarization
+To assess the effectiveness of ZEBRA binarization, we examined whether model performance can serve as an indicator of data quality.
+The results of this comparison are presented in table 1.
+Across all benchmark tasks, the Model Behavior-based scoring metric, SUP, demonstrates performance levels nearly equivalent to the instance-wise RLAIF method, with a minimal deviation of only 0.008.
+Notably, for MMLU-Pro and IFeval, ZEBRA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via Model Behavior can yield competitive or superior alignment outcomes.
+Table 1 shows the results of the average scores for each methodology.
+The detailed results, including average scores for each methodology, are summarized in Appendix E.
+Statistical significance.
+Table 2 reports paired -tests and permutation tests (10,000 shuffles) that compare each anchoring strategy with the baseline instance-wise RLAIF data across the same 1 024 evaluation prompts.
+None of the strategies achieves statistical significance at , although the hybrid SUP+SIM shows the largest absolute gain (, , ).
+These numbers indicate that ZEBRA’s zero-annotation binarization maintains baseline-level alignment quality (within 3 pp) without incurring per-instance scoring cost, even if a decisive improvement is not yet observed.
+5.3 RQ3: Cost & Efficiency Analysis
+Table 3 contrasts the labeling cost of ZEBRA with canonical RLHF and RLAIF pipelines.
+The standing out points is below: Absolute cost gap.
+LLM-annotated RLAIF corpora lower this to $0.6–$4 K by outsourcing each pairwise judgment to GPT-4, yet they still purchase every label.
+ZEBRA, by reusing benchmark leaderboards, pays no marginal cost ($0) for preference construction.
+Relative efficiency.
+Normalised per comparison, GPT-4 labels ($0.063) are cheaper than human labels ($0.67), but ZEBRA is orders of magnitude cheaper than both because it dispenses with pairwise annotation altogether.
+Cost matters only if quality survives. Despite a zero-dollar label budget, ZEBRA matches or surpasses RLAIF baseline (Table 1); the mean performance difference across six benchmarks is .
+Consequently, ZEBRA offers a cost-minimal, scalable, and annotation-free route to high-quality preference data.
+Because the price of human labor or GPT-4 tokens scales linearly with data volume, traditional pipelines become progressively more expensive as models, tasks, and safety domains proliferate.
+ZEBRA decouples alignment from annotation cost: adding a new model needs no further labels, and incorporating an extra benchmark simply augments the behavior matrix.
 6 Conclusion
-This paper introduces KARMA, a kinship-based framework for generating high-quality preference data for language model alignment without requiring instance-wise annotation.
-KARMA achieves performance comparable to RLAIF while offering greater computational efficiency and surpassing baselines on MMLU-Pro and IFeval, demonstrating the effectiveness of kinship-aware preference mapping.
-Our findings show that model size significantly affects the optimal kinship strategy.
-Smaller models (3B) benefit from prioritizing response quality, while larger models (7-8B) perform better when leveraging response similarity.
-Additionally, kinship patterns vary across tasks, with reasoning clustering by model size, instruction-following by model family, and knowledge-based tasks exhibiting a hierarchical structure.
-These results highlight the need for size-adaptive preference learning strategies and demonstrate that high-quality preference data can be generated without costly manual annotation.
-Future work will extend KARMA to cross-family model comparisons and further explore the relationship between model kinship and human preference judgments.
+We presented ZEBRA, a zero-annotation framework that replaces per-instance preference labeling with model-behavior knowledge distilled from public benchmark tables.
+By pairing responses according to either behavioral superiority or similarity—or a hybrid of the two—ZEBRA constructs 64k high-quality preference pairs at zero marginal cost, completely eliminating the dominant expense in RLHF or RLAIF pipelines.
+Experiments across 6 standardized benchmarks show that models fine-tuned with SUP achieve performance on par with, and sometimes better than, an RLAIF baseline.
+SIM and the hybrid SUP+SIM remain within a small, statistically insignificant margin of the baseline.
+These results confirm that benchmark-derived behavioral knowledge can serve as an effective proxy for preference supervision.
+Crucially, ZEBRA reduces marginal labeling cost to zero, reducing tens of thousands of dollars otherwise required for large-scale preference annotation.
+Future work will explore richer behavioral axes such as safety, toxicity, and multilingual ability; investigate adaptive schedules that incorporate additional refined pairing strategies during training; and evaluate how well ZEBRA preferences correlate with explicit human judgments.
 7 Related Work and Background
@@
 (2024).
-Various preference-based learning techniques, particularly reinforcement learning methods such as Proximal Policy Optimization (PPO)Schulman et al.
-(2017), Direct Preference Optimization (DPO)Rafailov et al.
-(2024), and Optimized Reinforcement Preference Optimization (ORPO)Hong et al.
+Various preference-based learning techniques, particularly reinforcement learning methodsSchulman et al.
+(2017); Rafailov et al.
+(2024); Hong et al.
 (2024), have been developed to facilitate this tuning process.
@@
 (2023) suite, which assess a model’s factual accuracy, reasoning ability, and compliance with instructions.
-While these benchmarks provide a foundation for selecting responses that promote alignment goals, existing methods typically overlook the impact of variation in model capabilities on alignment consistency.
-The KARMA Framework addresses this limitation by introducing a kinship-aware approach that emphasizes model compatibility in preference data selection.
-By curating preference pairs based on model kinship—aligning models with similar core competencies in knowledge, reasoning, and instruction-following—the KARMA Framework enhances alignment coherence and ensures stable, high-quality data.
-Further, its curriculum-based selection process sequences preference data from simpler to more complex distinctions, optimizing the alignment process for improved consistency and scalability.
+The ZEBRA Framework addresses this limitation by introducing a model behavior-level approach that emphasizes model compatibility in preference data selection.
+By curating preference pairs based on model behavior knowledge with similar core competencies in knowledge, reasoning, and instruction-following—the ZEBRA Framework enhances alignment coherence and ensures stable, high-quality data.
 Limitation
-In this paper, we demonstrated that KARMA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO.
+In this paper, we demonstrated that ZEBRA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO.
 However, due to resource and time constraints, we were unable to validate our approach across a broader range of alignment tuning techniques.
@@
 Additionally, our evaluation primarily focused on fundamental abilities such as Reasoning, Knowledge, and Instruction-Following.
-However, we did not assess KARMA’s performance on other important values, including factuality and ethical standard.
+However, we did not assess ZEBRA’s performance on other important values, including factuality and ethical standard.
 Future work should incorporate evaluations reflecting these aspects.
@@
 Appendix B Prompt template for Evaluation Benchmarks
-To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results.
+To determine behavior knowledge between models, this paper evaluates benchmark performance and compares the similarity of these numerical results.
 The benchmark performance of all 17 models considered in this study is presented in Table 7.
 For the start and end tokens, the template recommended in the paper was used.
-Appendix C Evaluation Origin Model Capabilities for Kinship
-To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results.
+Appendix C Evaluation Model Benchmark Performance
+To determine behavior knowledge relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results.
 The benchmark performance of all 17 models considered in this study is presented in Table 4.
@@
 Model IFeval MMLU-STEM MMLU-pro Hellaswag ARC-easy ARC-Challenge Average Score GPT-4 85.37 86.40 0.64 95.30 96.63 96.40 0.99 GPT-3.5-turbo 72.54 70.00 0.38 85.00 92.80 83.02 0.77 bard 89.33 71.80 0.59 84.70 84.43 77.13 0.86 Llama-2-7b-chat 25.19 53.10 0.20 67.50 72.14 54.61 0.41 Llama-2-13b-chat 24.82 57.80 0.25 71.20 72.05 58.02 0.45 Llama-2-70b-chat 24.07 68.90 0.38 78.10 82.20 67.66 0.58 UltraLM-13b 54.92 49.58 0.19 54.00 57.58 48.04 0.40 WizardLM-7b 45.83 42.50 0.18 77.70 39.48 34.90 0.34 WizardLM-13b 33.92 52.30 0.17 81.00 72.94 55.38 0.45 WizardLM-70b 49.51 52.70 0.27 83.30 80.68 71.93 0.58 Vicuna-33b 52.76 64.00 0.23 75.00 81.57 64.51 0.57 Alpaca-7b 30.58 37.92 0.15 23.60 43.31 33.79 0.16 Falcon-40b-instruct 24.54 67.50 0.14 80.00 76.60 56.70 0.47 MPT-30b-chat 30.70 50.40 0.20 24.53 87.12 70.73 0.38 Starchat 28.30 40.12 0.12 25.40 16.96 9.07 0.05 Pythia-12b 24.71 27.00 0.12 25.60 24.49 31.80 0.07
-Appendix D Total Performance
+Appendix D Benchmark Performance Representation
+What benchmark represent the model capability: knowledge, instruction-following, and reasoning?
+Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics.
+Notably, model behavior knowledge is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
+Analysis Framework
+To systematically investigate these relationships, we analyzed model similarities across three key capability dimensions:
+Knowledge-Based Tasks: Unlike reasoning capability, clustering in instruction-following tasks is more strongly aligned with model families rather than size.
+This indicates that training methodology and architectural choices exert a greater influence on instruction adherence.
+Reasoning Capability: Models tend to cluster primarily based on parameter count, suggesting that model size plays a dominant role in shaping reasoning performance.
+Instruction-Following (IF) Tasks: A hierarchical influence pattern emerges, where:
+At the initial hierarchy, the model family is the primary determinant.
+At the higher hierarchy, the model size becomes a stronger predictor of performance.
+Figure 11 visualizes these relationships through dendrograms, illustrating the hierarchical clustering patterns that emerge across different capability dimensions.
+Figure 8 shows the frequency of each model being selected as the positive or negative under each strategy. Overall, larger and more familiar models tend to be chosen more frequently.
+Appendix E Total Performance
 Table 5 provides an aggregated view of the overall performance of each model across all benchmark datasets.
 The total performance scores were computed by averaging the normalized scores across the selected evaluation metrics, offering a holistic comparison of model capabilities.
-| Benchmark | Model | Baseline | ||||
+Appendix F Model Size and ZEBRA Strategies
+The impact of different Strategies varies significantly with model size, influencing how models learn from preference data.
+To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters).
+Our analysis reveals a clear pattern in how model size determines the optimal strategy.
+Smaller models (3B) exhibit superior performance when trained using the SUP algorithm, which prioritizes learning from response quality within the binarized dataset.
+In contrast, larger models (7-8B) achieve better results with the SIM algorithm, suggesting that response similarity becomes increasingly important as model size grows.
+This trend indicates that model size fundamentally influences how different models leverage preference data.
+While smaller models benefit more from explicit learning based on absolute quality differences, larger models demonstrate greater sensitivity to the nuanced relationships between similar responses in the training data.
+This relationship is visualized in Figure 9, illustrating the distinct learning behaviors observed across different model scales.
+| Benchmark | Model | Baseline | SUP | SIM | SUP+SIM | |
 | MMLU-STEM | Llama-3.1-3B | SFT | 0.2331 | 0.2679 | 0.2434 | 0.2200 |